SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (23rd Jan, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Brandon Drummond, Al Eisner, Selina Li, Su Dong, Mark Tibbetts.

Operations

  • The boxes of old TSF boards have been moved from the EH to the trigger lab. The crate will be found a home in IR2.
  • The two colliding bemas runs taken on Friday (69572 and 69573) have been processed and the PR plots look fine for trigger data. Low luminosity runs like this ought to have the config key "PHYSICS_LOWLUMI" rather than "PHYSICS". This was mentionned in the 3.45 operations meeting.
  • While we're running cosmics still, it would be a good idea to swap out the spare BLT and GLT boards and check the PROMS are working properly on them.

EMT

  • The plots produced in PR are of L3 calibrated energy, rather than raw EMC energy (see here). An overlay would help to compare the changes between years of running. Also, the older plots include hot towers and the electronics problems the EMC used to have, so later plots would be a more direct comparison of changes in the M threshold (for example) over time. Check: what events are used to make the plots, how the axes are defined, what cuts applied etc.

L3

  • The new t-to-d functions have shown a very subtle effect, unless something unexpected show up they will be loaded (see hypernews discussion here and following).

Data Quality and Monitoring

  • Run 5 stripcharts will be posted shortly.

Config

  • Debbie showed some plots of single-track MC for electrons and muons, with (red) all tracks and (black) tracks from events where the 1Zk line fired. It's still not clear what they're showing, whether there's background included or whether PHOTOS is running in Moose. The same plots should be made for the 1Z trigger for comparison. Also, it appears impossible to acces the MC truth information in the event - speak to Eric about this.
  • Discussion of including the 1E trigger line - we could possibly put in a 1E line instead of the 1Z' line, by replacing the 3M/4 line. We ought to have a plan in case there's a real physics proposal soon.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.